# A distance function for stochastic matrices

### v1

@@ Abstract
Index Terms:
Information geometry, discrete Markov chain comparison, stochastic matrix comparison
I Introduction and background
Markov chains are a well studied tool in probability theory with numerous applications in communications [1, 2], simulation [3], queuing [4] and many other areas. Of particular interest to us is the use of Markov chains to describe operational processes [5, 6], and in particular those applicable to general healthcare services [7, 8, 9]. Of importance to population healthcare management are mental health [10], community care [11], and cost-effectiveness of interventions [12] which drive many decisions made by governments and professionals. In healthcare common questions about operational processes could include: What are the typical sequence of events when someone enters a hospital in an emergency? Can we redesign an operational process to be more efficient? How can we compare operational processes to see which are similar or different? In particular, the UK’s National Institute for Health Excellence (NICE) regularly publishes guidance and recommendations based on estimating transition probabilities which model healthcare processes as a Markov model [13]. It is this focus on comparing descriptions of health services, represented as Markov type models, which motivates us here [14, 15, 16, 17].
Within the domain of discrete time Markov chains, much work has been done on quantifying the similarity between two chains. The most widely used approach to comparing Markov chains is to consider the probabilities induced on the sequence space they generate, i.e comparing fixed length sequences of two models to distinguish different behaviour. Early work in this area [18] established that vectorised (hidden) Markov model transitions could be compared using simple functions: the dot product, cosine angle and Frobenius distance. Later, the use of the total variation distance on the space of traces generated by models was introduced [19], followed by the formulation of “linear” distances between samples from Markov chains [20]. A general approach to the $L^{p}$ norm–induced distances and the Kullback-Leibler divergence has also been studied [21], extending them to infinite time runs while exploring their properties of continuity.
@@ I Introduction and background
Although a step in a good direction, distances based on Markov sequences can in some scenarios still be overly sensitive to initial distributions or small changes in the stochastic matrix which represents a Markov model. For example a simple model of a biased coin toss, where outcome bias is parameterised by some real variable, can generate completely difference trace sequences if comparing two different initial distributions. This is obvious, given two different starting conditions we end up at different places. However the underlying generative process, as specified by the stochastic matrix, is the same. In keeping with healthcare, another example of where the initial state is of less importance is for hospital emergency department patient attendances or other specific disease pathways. As the process typically starts in the same initial state, e.g. “presented at hospital”, every comparison of processes across different organisations will use the same initial state distribution for a Markov chain. It therefore seems reasonable to separate the initial conditions from the Markov model’s structural properties. This amounts to equipping a distance function to the space of stochastic matrices. This suggests an approach to comparing Markov models via the structure of the stochastic matrices themselves, rather than traces generated by short or long term runs.
@@
Learning a stochastic matrix, given observation data, is also important in the fields of sensor-based human activity recognition [29] and designing statistically consistent algorithms for noisy data labelling [30]. Underlying these estimation techniques is the choice of objective function one might want to optimise. Typical objective functions used are: the direct calculation of the Kullback-Leibler divergence [31, 32] which is the usual divergence calculated row-wise and summed for a pair of stochastic matrices, some form of log-likelihood function as proposed for denoising in image analysis [33], the related cross entropy for video analysis [34], cross entropy hypothesis testing an estimated stochastic matrix against a reference [35] or even Bayesian based test approaches [36, 37]. In the majority of these settings short term runs are more important as recommendations for the next action are all that is needed, or the long time run of a Markov chain gives no more benefit after a finite number of steps, and the information contained within the stochastic matrix is all that is needed. Our observation is a well motivated comparison between Markov chain stochastic matrices would be beneficial, and could make more sense in certain scenarios than comparing sequences directly. This leads us to identify a stochastic matrix only distance function which mirrors desired properties of distances on sequence spaces.
@@
Throughout, we denote a time homogeneous discrete time Markov model with $M$ which is defined over a state space $X=\{x_{1},\ldots,x_{k}\}$. In other words we consider sequences of random variables $(Z_{1},Z_{2},\ldots)$ which satisfy the standard Markov property. Initial Markov chain distributions are written $\boldsymbol{\pi}$, with components $\pi(x_{i})$ representing the probability of being in state $x_{j}$, and the associated model’s (row) stochastic matrix $\boldsymbol{P}$, with components $P(x_{i},x_{j})=\mathrm{Pr}(Z_{t+1}=x_{j}|Z_{t}=x_{i})$. Sequences of length $n$ form a space $X^{n}$ and the space of infinite length sequences is written $X^{\omega}$. Over all elements of either $X^{n}$, or alternatively $X^{\omega}$, we can identify an induced probability vector $\boldsymbol{p}$, with components denoted $p(w)$, assigned for each sequence $w=(x_{1},\ldots,x_{n})\in X^{n}$, according to initial distribution $\boldsymbol{\pi}$ and stochastic matrix $\boldsymbol{P}$. Finally we make use of the Hadamard product (i.e. element wise product) between two elements, denoted as $\circ$, and the Hadamard square root (i.e. element wise square root) denoted as $\circ\frac{1}{2}$.
@@ II Sequence distances
To whit, we view the parameter space of a given statistical model as a manifold. The fundamental object in both information and differential geometry is the metric tensor. In simple terms, manifolds allow us to discuss concepts such as continuity and differentiation. A metric tensor allows us to define a notion of distance between points on a manifold, which means we can equip the parameter space of a statistical model with a distance function.
@@
There are many ways to define a metric tensor for a manifold and historically in information geometry this has been via what is known as the Fisher information metric tensor. Consider a parametric family of stochastic models whose distribution function is denoted $h(x;\boldsymbol{p})$. Here $x$ is an element of the probability’s support space $X$ and $\boldsymbol{p}\in\mathbb{R}^{n}$ is a vector of the parameters for the family, which we will shortly identify as a probability vector itself. Up to a constant scalar, the Fisher information metric at a point $\boldsymbol{p}$ is defined as an $n\times n$ matrix with components,
$g_{ij}(\boldsymbol{p})=-\mathbb{E}_{h}\left[\frac{\partial^{2}\log{h(x;\boldsymbol{p})}}{\partial p_{i}\partial p_{j}}\right]$ (1)
@@
The $g_{ij}$ are known as metric tensor components, and define the geometry of the parameter space, also known as a statistical manifold. We want to discuss the concept of geodesic curves, i.e. curves connecting two points of the parameter space which respect the underlying geometry of the space. Geodesics are derived using the geodesic equation which in turn is derived from the metric tensor. Intuitively, the solutions to the geodesic equations are the “shortest” path curves with respect to the geometry of the space under consideration. Conventionally, geodesic curves are parameterised by some real variable $s$, which encodes the start point and end point of the curve. Ultimately we want to find a closed form for the induced distance function, minimised over all possible geodesic curves $\boldsymbol{c}(s)$. In the case of a parametric family of probability distributions these points are denoted $\boldsymbol{c}(s_{1})=\boldsymbol{p}_{1}$ and $\boldsymbol{c}(s_{2})=\boldsymbol{p}_{2}$. In general, there can be multiple geodesic curves connecting two points. For example the two ways round a great circle on a sphere. Thus we define the distance between two points using the geodesic curve which has the smallest length,
$d(\boldsymbol{p}_{1},\boldsymbol{p}_{2})=\min_{\boldsymbol{c}(s)}\int_{s_{1}}^{s_{2}}\mathrm{d}s\sqrt{\sum_{i,j}g_{ij}(\boldsymbol{p})\frac{\mathrm{d}c_{i}(s)}{\mathrm{d}s}\frac{\mathrm{d}c_{j}(s)}{\mathrm{d}s}}$ (2)
@@
$d(\boldsymbol{p}_{1},\boldsymbol{p}_{2})$ would then be our desired distance function on a statistical manifold. It is well known that Markov chains induce a probability distribution over sequences of fixed length [39]. These distributions correspond to categorical distributions and are defined as having the probability mass function,
$h(x;\boldsymbol{p})=\prod_{i}p(x_{i})^{[x=x_{i}]}$ (3)
@@
with domain $X=\{x_{1},\ldots,x_{n}\}$ and $\boldsymbol{p}=(p(x_{1}),\ldots,p(x_{n}))$. In other words the probability of observing each possible event $x_{i}$ is $p(x_{i})$, akin to the probability distribution of rolling an $n$-sided dice. For a fixed state space of $n$ elements, the space of all probabilities which can parameterise a categorical distribution is known as the simplex space $\Delta=\{\boldsymbol{p}:\|\boldsymbol{p}\|_{1}=1\}$. Thus from here we can compute the metric tensor components in Eq. 1 and work through until we arrive at a solution for Eq. 2.
@@
$g_{ij}(\boldsymbol{p})=\frac{\delta_{ij}}{p_{i}}$ (4)
Remarkably, we can sidestep the usual geodesic equation challenge associated with differential geometry. The coordinate transformation $\boldsymbol{y}=\boldsymbol{p}^{\circ\frac{1}{2}}$ allows us to show the metric tensor components become $g_{ij}(\boldsymbol{y})=\delta_{ij}$. This transformation shows the components are the standard Euclidean metric in Euclidean space, with the condition that points are constrained to a positive quadrant of unit hypersphere centered at the origin ($\boldsymbol{y}\cdot\boldsymbol{y}=1)$. Hence, the associated geodesic curves we want are the great arcs between points $\boldsymbol{y}_{1}$ and $\boldsymbol{y}_{2}$. As a result, a closed form for the distance between two categorical distributions with $n$ states and parameterised with $\boldsymbol{p}_{1},\boldsymbol{p}_{2}\in\Delta$ is,
$d^{(n)}(\boldsymbol{p}_{1},\boldsymbol{p}_{2})=2\arccos\mathrm{BC}(\boldsymbol{p}_{1},\boldsymbol{p}_{2})$ (5)
@@
$\mathrm{BC}(\boldsymbol{p}_{1},\boldsymbol{p}_{2})=\sum_{i=1}^{n}\sqrt{p_{1}(x_{i})p_{2}(x_{i})}$ (6)
@@
The final link back to Markov chains comes from realising that the collection of all probabilities of observing an $n$ step Markov chain trace forms a categorical distribution over $X^{n}$. Explicitly, denoting a general sequence of events as $w=(x_{1},x_{2},\ldots,x_{n})$ for a chain $M$ and $n\geq 2$, the probability of observing the trace $w$ given an initial distribution $\boldsymbol{\pi}$ for a Markov chain with stochastic matrix $\boldsymbol{P}$ is,
@@
Thus we arrive at a distance for Markov chains,
@@
The key concepts to remember here are: 1) using information geometry, a notion of distance can be naturally motivated for Markov chains, 2) the distance satisfies all properties of a true distance function and 3) is fundamentally built to respect the basic probability properties of a Markov chain. Following [45], for convenience we introduce the following definitions, where $\circ$ denote Hadamard operations,
$\boldsymbol{r}$ $=\boldsymbol{\pi}_{1}^{\circ\frac{1}{2}}\circ\boldsymbol{\pi}_{2}^{\circ\frac{1}{2}}$ (9a)
$\boldsymbol{R}$ $=\boldsymbol{P}_{1}^{\circ\frac{1}{2}}\circ\boldsymbol{P}_{2}^{\circ\frac{1}{2}}$ (9b)
Considering the product structure for each $p_{i}(w)$, we can compactly write the sum over length $n$ words,
$d^{(n)}(M_{1},M_{2})=2\arccos\left\|\boldsymbol{r}^{T}\boldsymbol{R}^{n}\right\|_{1}$ (10)
@@
Note this is equivalent to the matrix product form of $d^{\omega}(M_{1},M_{2})=\lim_{n\rightarrow\infty}2\arccos\left\|\boldsymbol{r}^{T}\boldsymbol{R}^{n}\right\|_{1}$. We can see the Bhattacharyya angle always exists in the limit $n\rightarrow\infty$ by observing (via the inequality of arithmetic and geometric means) $\sqrt{p_{1}(w)p_{2}(w)}\leq\frac{p_{1}(w)+p_{2}(w)}{2}$ and using the direct comparison test for series convergence [47]. As $\sum_{w\in X^{\omega}}\frac{p_{1}(w)+p_{2}(w)}{2}=1$ (i.e. converges) the direct comparison test proves $d^{\omega}(M_{1},M_{2})$ converges. Convergence of the Bhattacharyya angle is then ensured by the continuity of $\arccos$ on $[0,1]$.
A simple characterisation of when $d^{\omega}(M_{1},M_{2})$ is guaranteed to attain its maximum can be found from the inequality $\rho(\boldsymbol{P}_{1}^{\circ\frac{1}{2}}\circ\boldsymbol{P}_{2}^{\circ\frac{1}{2}})\leq\rho(\boldsymbol{P}_{1})\rho(\boldsymbol{P}_{2})=1$ (see Huang et. al. for details [48]). In other words when the spectral radius of $\boldsymbol{P}_{1}^{\circ\frac{1}{2}}\circ\boldsymbol{P}_{2}^{\circ\frac{1}{2}}$ is strictly less than unity the distance converges to the numerical constant $\pi$ in the limit $n\rightarrow\infty$. This means the long term run of two such Markov models are statistically distinguishable according to the Bhattacharyya angle. We can however go further and study two types of matrices for which a closed form for the “Bhattacharyya rate” $d^{\omega}(M_{1},M_{2})$ is immediately computable and not trivially a maximum [45],
@@
    Type 1: Diagonalisable matrices (over $\mathbb{C}$)
@@
Primitive matrices are defined as non-negative matrices (i.e. all elements are non-negative) such that for some positive integer $k>0$ the matrix power $\boldsymbol{R}^{k}$ is a strictly positive matrix (i.e. all elements are positive). Primitive matrices play a central role in various proofs based on the Perron-Frobenius theorem [49].
Although with substantial overlap, these matrix types do have non-trivial complements in the sense that there are matrices which exist as one type and not the other. Denoting the eigenvalues, right eigenvectors and left eigenvectors of $\boldsymbol{R}$ as $\lambda_{i}$, $\boldsymbol{p}_{i}$ and $\boldsymbol{q}_{i}$ respectively, the results we need are,
$\lim_{n\rightarrow\infty}\boldsymbol{R}^{n}=\begin{cases}\sum_{i\in\mathcal{I}}\boldsymbol{p}_{i}\boldsymbol{q}_{i}^{T}&\text{: type 1}\\ \boldsymbol{p}_{1}\boldsymbol{q}_{1}^{T}&\text{: type 2}\end{cases}$ (12)
With the associated requirements being,
@@
    Type 1: $\boldsymbol{R}$ is diagonalisable, $\rho(\boldsymbol{R})=1$, $-1\notin\mathrm{spec}(\boldsymbol{R})$ and =$\mathcal{I}=\{i:\lambda_{i}=1,\lambda_{i}\in\mathrm{spec}(\boldsymbol{R})\}$
@@
    Type 2: $\boldsymbol{R}$ is primitive, $\rho(\boldsymbol{R})=\lambda_{1}=1$
Inserting the above into $d^{\omega}(M_{1},M_{2})$ we can collect the various results as follows,
$d^{\omega}(M_{1}M_{2})=\begin{cases}2\arccos\sum\limits_{i\in\mathcal{I}}\|\boldsymbol{r}^{T}\boldsymbol{p}_{i}\boldsymbol{q}_{i}^{T}\|_{1}&\text{: type 1}\\ 2\arccos\|\boldsymbol{r}^{T}\boldsymbol{p}_{1}\boldsymbol{q}_{1}^{T}\|_{1}&\text{: type 2}\\ \pi&\text{: $\rho(\boldsymbol{R})<1$}\end{cases}$ (13)
@@
Markov chain mixing times are concerned with estimating a finite time step $\tau_{\star}$ such that the distance between some reference stationary distribution and the $\tau_{\star}$ step distribution of some Markov chain is “close”. Denoting an initial Markov chain distribution for a given state $x_{j}\in X$ as $\boldsymbol{e}_{j}$ (i.e. the standard basis in $\mathbb{R}^{n}$), we define the Bhattacharyya angle induced mixing time of a Markov chain $\boldsymbol{P}$ (with dimensions $n\times n$) starting in state $\boldsymbol{e}_{j}$, with unique stationary distribution $\boldsymbol{\pi}_{\star}$, as
$\tau_{\star}(x)=\min\left\{\tau>0:d^{(n)}\left(\boldsymbol{\pi}_{\star},\boldsymbol{e}_{j}^{T}\boldsymbol{P}^{\tau}\right)\leq\epsilon\right\}$ (14)
We can operationally call two Markov chains “similar’ if after some time $\tau_{\star}$ their state distributions are “close enough” for the problem at hand. Another way to view this is that if we have a obtained a mixing time for one chain, how can we use it to bound the mixing time of another chain? In some sense one can then approximate one Markov chain with another after a finite number of steps, when comparing the marginal state distributions of two Markov chains.
A common class of Markov chains considered in mixing times are ergodic (irreducible and aperiodic) and reversible, and share the property of having a single unique stationary distribution. Under these assumptions, bounds on $\tau_{\star}$ can be placed as we can take advantage of the fact the $\tau$-step transition matrix components can be written as [51],

### v2

@@ Abstract
preprint: APS/123-QED
I Introduction
Markov chains are a well studied tool in probability theory with numerous applications in communications [1, 2], simulation [3], queuing [4] and many other areas. Of particular interest to us is the use of Markov chains to describe operational processes [5, 6], and in particular those applicable to general healthcare services [7, 8, 9]. Of importance to population healthcare management are mental health [10], community care [11], and cost-effectiveness of interventions [12] which drive many decisions made by governments and professionals. In healthcare common questions about operational processes could include: What are the typical sequence of events when someone enters a hospital in an emergency? Can we redesign an operational process to be more efficient? How can we compare operational processes to see which are similar or different? In particular, the UK’s National Institute for Health Excellence (NICE) regularly publishes guidance and recommendations based on estimating transition probabilities which model healthcare processes as a Markov chain [13], and for example see [14] for an explicit example of a transition matrix defined using clinical test results with respect to Peginterferon beta-1a for treating relapsing–remitting multiple sclerosis. It is this focus on comparing descriptions of health services, represented as Markov type models, which motivates us here [15, 16, 17, 18].
Within the domain of discrete time Markov chains, much work has been done on quantifying the similarity between two chains. The most widely used approach to comparing Markov chains is to consider the probabilities induced on the sequence space they generate, i.e comparing fixed length sequences of two models to distinguish different behaviour. Early work in this area [19] established that vectorised (hidden) Markov chain transitions could be compared using simple functions: the dot product, cosine angle and Frobenius distance. Later, the use of the total variation distance on the space of traces generated by models was introduced [20], followed by the formulation of “linear” distances between samples from Markov chains [21]. A general approach to the $L^{p}$ norm–induced distances and the Kullback-Leibler divergence has also been studied [22], extending them to infinite time runs while exploring their properties of continuity.
@@ I Introduction
Although a step in a good direction, distances based on Markov sequences can in some scenarios still be overly sensitive to initial distributions or small changes in the stochastic matrix which represents a Markov chain. For example a simple model of a biased coin toss, where outcome bias is parameterised by some value $p\in[0,1]$, can generate completely difference trace sequences if comparing two different initial distributions. This is obvious, given two different starting conditions we end up at different places. However the underlying generative process, as specified by the stochastic matrix, is the same. In keeping with healthcare, another example of where the initial state is of less importance is for hospital emergency department patient attendances or other specific disease pathways. As the process typically starts in the same initial state, e.g. “presented at hospital”, every comparison of processes across different organisations will use the same initial state distribution for a Markov chain. It therefore seems reasonable to separate the initial conditions from the Markov chain’s structural properties. This amounts to equipping the space of stochastic matrices with a distance function. This suggests an approach to comparing Markov chains via the structure of the stochastic matrices themselves, rather than traces generated by short or long term runs.
@@
Learning a stochastic matrix, given observation data, is also important in the fields of sensor-based human activity recognition [30] and designing statistically consistent algorithms for noisy data labelling [31]. Underlying these estimation techniques is the choice of objective function one might want to optimise. Typical objective functions used are: the direct calculation of the Kullback-Leibler divergence [32, 33] which is the usual divergence calculated row-wise and summed for a pair of stochastic matrices, some form of log-likelihood function as proposed for denoising in image analysis [34], the related cross entropy for video analysis [35], cross entropy hypothesis testing an estimated stochastic matrix against a reference [36] or even Bayesian based test approaches [37, 38]. In the majority of these settings short term runs are more important as recommendations for the next action are all that is needed, or the long time run of a Markov chain gives no more benefit after a finite number of steps, and the information contained within the stochastic matrix is all that is needed. Our observation is that a well motivated comparison between Markov chain stochastic matrices would be beneficial, and could make more sense in certain scenarios than comparing sequences directly. This leads us to identify a distance function between stochastic matrices which mirrors desired properties of distances on sequence spaces.
@@
Throughout, we denote a time homogeneous discrete time Markov chain with $M$ which is defined over a state space $X=\{x_{1},\ldots,x_{k}\}$. In other words we consider sequences of random variables $(Z_{1},Z_{2},\ldots)$ which satisfy the standard Markov property. Initial Markov chain distributions are written $\bm{\pi}$, with components $\pi(x_{i})$ representing the probability of being in state $x_{j}$, and the associated model’s (row) stochastic matrix $\bm{P}$, with components $P(x_{i},x_{j})=\mathrm{Pr}(Z_{t+1}=x_{j}|Z_{t}=x_{i})$. Sequences of length $n$ form a space $X^{n}$ and the space of infinite length sequences is written $X^{\omega}$. Over all elements of either $X^{n}$, or alternatively $X^{\omega}$, we can identify an induced probability vector $\bm{p}$, with components denoted $p(w)$, assigned for each sequence $w=(x_{1},\ldots,x_{n})\in X^{n}$, according to initial distribution $\bm{\pi}$ and stochastic matrix $\bm{P}$. The Iverson bracket is written as $[A]$ for some boolean statement $A$, and the collection of eigenvalues of a matrix $\bm{R}$ is denoted $\sigma(\bm{R})$. Finally we make use of the Hadamard product (i.e. element wise product) between two elements, denoted as $\circ$, and the Hadamard square root (i.e. element wise square root) denoted as $\circ\frac{1}{2}$.
@@ II Sequence distances
We treat the parameter space of a given statistical model as a manifold. The fundamental object in both information and differential geometry is the metric tensor. Manifolds allow us to discuss concepts such as continuity and differentiation. A metric tensor allows us to define a notion of distance between points on a manifold, which means we can equip the parameter space of a statistical model with a distance function.
@@
There are many ways to define a metric tensor for a manifold and historically in information geometry this has been via the Fisher information metric tensor. Consider a parametric family of stochastic models whose distribution function is denoted $h(x;\bm{p})$. Here $x$ is an element of the distribution’s support space $X$ and $\bm{p}\in\mathbb{R}^{n}$ is a vector of the parameters for the family, which we will show can be treated as a probability itself. Up to a constant scalar, the Fisher information metric at a point $\bm{p}$ is defined as an $n\times n$ matrix with components,
$g_{ij}(\bm{p})=-\mathbb{E}_{h}\left[\frac{\partial^{2}\log{h(x;\bm{p})}}{\partial p_{i}\partial p_{j}}\right]$ (1)
@@
where $\mathbb{E}_{h}$ indicates the expectation value is taken with respect to the probability distribution $h(x;\bm{p})$ over the support of $x$. The $g_{ij}$ are known as metric tensor components, and define the geometry of the parameter space, which is a statistical manifold. We want to discuss the concept of geodesic curves, i.e. curves connecting two points of the parameter space which respect the underlying geometry of the space. Geodesics are derived using the geodesic equation which in turn is derived from the metric tensor. Intuitively, the solutions to the geodesic equations are the “shortest” path curves with respect to the geometry of the space under consideration. Conventionally, geodesic curves are parameterised by some real variable $s$, which encodes the start point and end point of the curve. Ultimately we want to find a closed form for the induced distance function, minimised over all possible geodesic curves $\bm{c}(s)$. In the case of a parametric family of probability distributions these points are denoted $\bm{c}(s_{1})=\bm{p}_{1}$ and $\bm{c}(s_{2})=\bm{p}_{2}$. In general, there can be multiple geodesic curves connecting two points. For example the two ways round a great circle on a sphere. Thus we define the distance between two points using the geodesic curve which has the smallest length,
$d(\bm{p}_{1},\bm{p}_{2})=\min_{\bm{c}(s)}\int_{s_{1}}^{s_{2}}\mathrm{d}s\sqrt{\sum_{i,j}g_{ij}(\bm{p})\frac{\mathrm{d}c_{i}(s)}{\mathrm{d}s}\frac{\mathrm{d}c_{j}(s)}{\mathrm{d}s}}$ (2)
@@
$d(\bm{p}_{1},\bm{p}_{2})$ would then be our desired distance function on a statistical manifold. It is well known that Markov chains induce a probability distribution over sequences of fixed length [40]. These distributions correspond to categorical distributions and are defined as having the probability mass function,
$h(x;\bm{p})=\prod_{i}p(x_{i})^{[x=x_{i}]}$ (3)
@@
with domain of input events $X=\{x_{1},\ldots,x_{n}\}$ and associated probabilities $\bm{p}=(p(x_{1}),\ldots,p(x_{n}))$. In other words the probability of observing each possible event $x_{i}$ is $p(x_{i})$, akin to the probability distribution of rolling an $n$-sided dice. For a fixed state space of $n$ elements, the space of all probabilities which can parametrise a categorical distribution is known as the simplex space $\Delta=\{\bm{p}:\|\bm{p}\|_{1}=1\}$. Thus from here we can compute the metric tensor components in Eq. 1 and work through until we arrive at a solution for Eq. 2.
@@
$g_{ij}(\bm{p})=\frac{\delta_{ij}}{p_{i}}$ (4)
where $\delta_{ij}$ is a Kronecker delta. In differential geometry a metric tensor can be used to define geodesic equations, i.e. equations which define in some sense the “straightest” path. Typically solving these equations is difficult, however in this case we can sidestep the usual challenge. The coordinate transformation $\bm{y}=\bm{p}^{\circ\frac{1}{2}}$ allows us to show the metric tensor components become $g_{ij}(\bm{y})=\delta_{ij}$. This transformation shows the components are the standard Euclidean metric in Euclidean space, with the condition that points are constrained to a positive quadrant of unit hypersphere centered at the origin ($\bm{y}\cdot\bm{y}=1)$. Hence, the associated geodesic curves we want are arcs following the great circles defined by points $\bm{y}_{1}$ and $\bm{y}_{2}$. As a result, a closed form for the distance between two categorical distributions with $n$ states and parameterised with $\bm{p}_{1},\bm{p}_{2}\in\Delta$ is,
$d^{(n)}(\bm{p}_{1},\bm{p}_{2})=2\arccos\mathrm{BC}(\bm{p}_{1},\bm{p}_{2})$ (5)
@@
$\mathrm{BC}(\bm{p}_{1},\bm{p}_{2})=\sum_{i=1}^{n}\sqrt{p_{1}(x_{i})p_{2}(x_{i})}$ (6)
@@
The final link back to Markov chains comes from realising that the collection of all probabilities of observing an $n$ step Markov chain trace forms a categorical distribution over $X^{n}$. Explicitly, denoting a general sequence of events as $w=(x_{1},x_{2},\ldots,x_{n})$ for a chain $M$ and $n\geq 2$, the probability of observing the trace $w$ given an initial distribution $\bm{\pi}$ for a Markov chain with stochastic matrix $\bm{P}$ is,
@@
Making the identification $M_{1}\rightarrow\bm{p}_{1}$, $M_{2}\rightarrow\bm{p}_{2}$ we arrive at a distance for Markov chains,
@@
The key concepts here are: 1) using information geometry, a notion of distance can be naturally motivated for Markov chains, 2) the distance satisfies all properties of a true distance function and 3) the distance is constructed to respect the basic probability properties of a Markov chain. Following [46], for convenience we introduce the following definitions, where $\circ$ denote Hadamard operations,
$\bm{r}$ $=\bm{\pi}_{1}^{\circ\frac{1}{2}}\circ\bm{\pi}_{2}^{\circ\frac{1}{2}}$ (9a)
$\bm{R}$ $=\bm{P}_{1}^{\circ\frac{1}{2}}\circ\bm{P}_{2}^{\circ\frac{1}{2}}$ (9b)
Considering the product structure for each $p_{i}(w)$ and denoting the vector of all ones as 1, we can compactly write the sum over length $n$ words,
$d^{(n)}(M_{1},M_{2})=2\arccos\left(\bm{r}^{T}\bm{R}^{n}\textbf{1}\right)$ (10)
@@
Note this is equivalent to the matrix product form of $d^{\omega}(M_{1},M_{2})=\lim_{n\rightarrow\infty}2\arccos\left(\bm{r}^{T}\bm{R}^{n}\textbf{1}\right)$. We can see the Bhattacharyya angle always exists in the limit $n\rightarrow\infty$ by observing (via the inequality of arithmetic and geometric means) $\sqrt{p_{1}(w)p_{2}(w)}\leq\frac{p_{1}(w)+p_{2}(w)}{2}$ and using the direct comparison test for series convergence [48]. As $\sum_{w\in X^{\omega}}\frac{p_{1}(w)+p_{2}(w)}{2}=1$ (i.e. converges) the direct comparison test proves $d^{\omega}(M_{1},M_{2})$ converges. Convergence of the Bhattacharyya angle is then ensured by the continuity of $\arccos$ on $[0,1]$.
A simple characterisation of when $d^{\omega}(M_{1},M_{2})$ is guaranteed to attain its maximum can be found from the inequality $\rho(\bm{P}_{1}^{\circ\frac{1}{2}}\circ\bm{P}_{2}^{\circ\frac{1}{2}})\leq\rho(\bm{P}_{1})\rho(\bm{P}_{2})=1$ (see Huang et. al. for details [49]). In other words when the spectral radius of $\bm{P}_{1}^{\circ\frac{1}{2}}\circ\bm{P}_{2}^{\circ\frac{1}{2}}$ is strictly less than unity the distance converges to the numerical constant $\pi$ in the limit $n\rightarrow\infty$. This means the long term run of two such Markov chains are statistically distinguishable according to the Bhattacharyya angle. We can however go further and study two types of matrices for which a closed form for the “Bhattacharyya rate” $d^{\omega}(M_{1},M_{2})$ is immediately computable and not trivially a maximum [46],
@@
    Type 1: Diagonalisable matrices (over $\mathbb{C}$) with real eigenvalues
@@
Primitive matrices are defined as non-negative matrices (i.e. all elements are non-negative) such that for some positive integer $k>0$ the matrix power $\bm{R}^{k}$ is a strictly positive matrix (i.e. all elements are positive). Primitive matrices play a central role in various proofs based on the Perron-Frobenius theorem [50]. Although with substantial overlap, these matrix types do have non-trivial complements in the sense that there are matrices which exist as one type and not the other.
Theorem 1 (Bhattacharyya rate).
The Bhattacharyya rate between two Markov chain models which satisfy type 1 or type 2 requirements is
$d^{\omega}(M_{1},M_{2})=\begin{cases}2\arccos\sum\limits_{i\in\mathcal{I}}\bm{r}^{T}\bm{p}_{i}\bm{q}_{i}^{T}\bm{1}&\mathrm{:type~{}1}\\ 2\arccos\left(\bm{r}^{T}\bm{p}_{1}\bm{q}_{1}^{T}\bm{1}\right)&\mathrm{:type~{}2}\\ \pi&\mathrm{:~{}}\rho(\bm{R})<1\end{cases}$ (12)
Proof.
Define the requirements of the two types of Markov chains as
@@
    Type 1: $\bm{R}$ is diagonalisable, $\rho(\bm{R})=1$, $\lambda_{i}>-1$, and =$\mathcal{I}=\{i:\lambda_{i}=1,\lambda_{i}\in\sigma(\bm{R})\}$
@@
    Type 2: $\bm{R}$ is primitive, $\rho(\bm{R})=\lambda_{1}=1$
Denote the eigenvalues, right eigenvectors and left eigenvectors of $\bm{R}$ as $\lambda_{i}$, $\bm{p}_{i}$ and $\bm{q}_{i}$ respectively. The definition of diagonalisation allows us to write type 1 matrices as $\bm{R}=\sum_{i}\lambda_{i}\bm{p}_{i}\bm{q}_{i}^{T}$. Thus we can write an arbitrary matrix power as $\bm{R}^{n}=\sum_{i}\lambda_{i}^{n}\bm{p}_{i}\bm{q}_{i}^{T}$. The eigenvalues with absolute value strictly less than one vanish in the limit $n\rightarrow\infty$. Eigenvalues equal to unity are the only ones which remain. In a similar way, the Perron-Frobenius theorem states any type 2 matrix (i.e. primitive matrices) has the limit $\bm{R}^{n}\rightarrow\bm{p}_{1}\bm{q}_{1}^{T}$ as $n\rightarrow\infty$. Therefore we can collect the results as
$\lim_{n\rightarrow\infty}\bm{R}^{n}=\begin{cases}\sum_{i\in\mathcal{I}}\bm{p}_{i}\bm{q}_{i}^{T}&\text{: type 1}\\ \bm{p}_{1}\bm{q}_{1}^{T}&\text{: type 2}\end{cases}$ (13)
Inserting the above into Eq. 10 we obtain a closed form for $d^{\omega}(M_{1},M_{2})$ and arrive at our result. ∎
@@
Markov chain mixing times are concerned with estimating a finite time step $\tau_{\star}$ such that the distance between some reference stationary distribution and the $\tau_{\star}$ step distribution of some Markov chain is “close”. Denoting an initial Markov chain distribution for a given state $x_{j}\in X$ as $\bm{e}_{j}$ (i.e. the standard basis in $\mathbb{R}^{n}$), we define the Bhattacharyya angle induced mixing time of a Markov chain $\bm{P}$ (with dimensions $n\times n$) starting in state $\bm{e}_{j}$, with unique stationary distribution $\bm{\pi}_{\star}$, as
$\tau_{\star}(x)=\min\left\{\tau>0:d^{(n)}\left(\bm{\pi}_{\star},\bm{e}_{j}^{T}\bm{P}^{\tau}\right)\leq\epsilon\right\}$ (14)
We can operationally call two Markov chains “similar” if after some time $\tau_{\star}$ their state distributions are “close enough” for the problem at hand. Another way to view this is that if we have a obtained a mixing time for one chain, how can we use it to bound the mixing time of another chain? In some sense one can then approximate one Markov chain with another after a finite number of steps, when comparing the marginal state distributions of two Markov chains.
A common class of Markov chains considered in mixing times are ergodic (irreducible and aperiodic) and reversible. Ergodic means that a discrete finite Markov chain is,
  1. 1.
    Irreducible: $\forall i,j\,\exists\,\tau\textrm{ such that }P^{(n)}_{ij}>0$, i.e. all states are reachable from another in a finite number of steps $\tau$
  2. 2.
    Aperiodic: The greatest common divisor of the lengths of all cycles of the Markov chain’s stochastic matrix is one
and a well known result is that ergodic chains have a unique stationary distribution $\bm{\pi}$. For an ergodic chain, reversible means the associated unique stationary distribution and stochastic matrix satisfy,
$\mathrm{diag}(\bm{\pi})\,\bm{P}=\bm{P}^{T}\,\mathrm{diag}(\bm{\pi})$ (15)
which are known as the detailed balance equations. Under these assumptions, bounds on $\tau_{\star}$ can be placed as we can take advantage of the fact that the $\tau$-step transition matrix components can be written as [52],